Find the Taylor series for f centered at 42n+1(26) = (-1) 132n+1 and 42m(26) = 0 for all n. Rx) = 8 n=0
Answer #1

Given penti) ( 7 ) = (-2 (1372nt o and pen (A)=0 ... by 0 Buz y=0, 1, 2, 3 ....... we have p (O+1) (+) - GDO (137071 for m=l78 () = 7°(E) + P*) (6-8) + (15) (E) + so (*3*+ **(FE) (x+*+. 776)= PLE) + P°CE5) 6x=75) + PLE) (x - + ?($)(n+8) + FCLF) (x=